In a round tournament every team plays every other team once.
Example: For a three teams tournament 3 games would be played.
(a) How many games are played if there are four teams?
(b) How many games are played if there are five teams?
(c) How many games are played if there are n teams?


Answer by stanbon(75887)

In a round tournament every team plays every other team once.
Example: For a three teams tournament 3 games would be played.
(a) How many games are played if there are four teams?
Ans: 4C2 = (4*3)/1*2) = 6
------------------------------------------
(b) How many games are played if there are five teams?
Ans: 5C2 = (5*4)/(1*2) = 10
-----------------------------------------
(c) How many games are played if there are n teams
Ans: nC2 = (n(n-1))/2
==========================
Cheers,
Stan H.